OpenStudy (dj3strella):

help please 1. Which pair of numbers is 6 a common factor of? (Points : 1) 12 and 20 16 and 24 24 and 54 30 and 64

OpenStudy (anonymous):

c

OpenStudy (dj3strella):

Yes

OpenStudy (dj3strella):

thank you

OpenStudy (anonymous):

6x4=24, 6x9=54

OpenStudy (anonymous):

Your Welcome :)

OpenStudy (timaashorty):

So 12= 6 *2 ; 20 = 6 * ? ( NOT THIS ONE) 16= 6* ? ( not this ONE) \[\huge 24= 6*4 ; 54 = 6 * 9 ( THIS \ ONE \ is \ correct ) \]

OpenStudy (timaashorty):

\[\huge the\ third\ one\]